Mia Benson invested her summer earnings of R3 000 in a savings account which pays 2.5% interest. How much will this amount be in 6 months?

1
Expert's answer
2021-11-23T13:28:09-0500

"FV=PV(1+r)^n"


after 6 months:

"FV=3000(1+0.025)^{1/2}=R\ 3037.27"